The cost function is given by÷

C(y)=120y-y^2+0.02^3.

Obtain the following :

1.The average variable cost

2.The marginal cost


1
Expert's answer
2021-02-10T15:21:39-0500

(a) The average variable cost can be found as follows:


"C(y)_{avg}=\dfrac{C(y)}{y}=\dfrac{120y-y^2+0.02^3}{y},""C(y)_{avg}=120-y+\dfrac{0.02^3}{y}."

(b)

"MC=\dfrac{d}{dy}(C(y)),""MC=\dfrac{d}{dy}(120y-y^2+0.02^3)=120-2y."